OpenStudy (emmynimmy):

Use the given information about the polynomial graph to write the equation. Degree 3. Zeros at x = −5, x = −2, and x = 1. y-intercept at (x, y) = (0, 9). so y equals what

OpenStudy (sooobored):

zeros at x=-5, x=-2, and x=1 means that x+5=0, x+2=0, and x-1=0 are possible solutions of your polynomial

OpenStudy (emmynimmy):

okay so what does y equal

OpenStudy (issimplcalcus):

You can start by multiplying the zeros (FOIL-ing), you'd write them as if you had factored them. (x + 5)(x + 2)(x - 1)

OpenStudy (sooobored):

finding the root is when you set y equal to zero and solve for x hence your polynomial should incorporate (x+5), (x+2), and (x-1) like calcus wrote above

OpenStudy (emmynimmy):

okay but can u solve it for me so i can see how u did it

OpenStudy (sooobored):

its not solved fyi

OpenStudy (emmynimmy):

okay so can one of you solve it.

OpenStudy (emmynimmy):

??

OpenStudy (sooobored):

looking at the information given, Degree 3 this means that the highest exponent in our polynomial will be 3 or x^3

OpenStudy (sooobored):

meaning that our polynomial will have at most only 3 solutions/roots

OpenStudy (sooobored):

this also means that the solution will approximately look like y~ (x+5)(x+2)(x-1) and that the roots given are all singular roots as in, it wouldnt look like \[(x+5)^2(x+2)^4 (x-1)\]

OpenStudy (sooobored):

the y-intercept is given as (0,9) this means when you substitute x=0. then y should be equal to 9 however, what value do you get if you substitute x=0 in the expression below? (x+5)(x+2)(x-1)